Great Northwoods Outfitters is a retail phone-catalog company that specializes in outdoor clothing and equipment. A phone station at the company will be staffed with either full-time operators or temporary operators 8 hours per day. Full-time operators, because of their experience and training, process more orders and make fewer mistakes than temporary operators. However, temporary operators are cheaper because they receive a lower wage rate and are not paid benefits. A full-time operator can process about 360 orders per week, whereas a temporary operator can process about 270 orders per week. A full-time operator averages 1.1 defective orders per week, and a part-time operator incurs about 2.7 defective orders per week. The company wants to limit defective orders to 200 per week. The cost of staffing a station with full-time operators is $610 per week, and the cost of a station with part-time operators is $450 per week. Using historical data and forecasting techniques, the company has developed estimates of phone orders for an 8-week period, as follows:

Week Orders Week Orders
1 19,500 5 33,400
2 21,000 6 29,800
3 25,600 7 27,000
4 27,200 8 31,000

The company does not want to hire or dismiss full-time employees after the first week (i.e., the company wants a constant group of full-time operators over the 8-week period). The company wants to determine how many full-time operators it needs and how many temporary operators to hire each week to meet weekly demand while minimizing labor costs.
a. Formulate a linear programming model for this problem
b. Solve this model by using the computer (using Excel Solver)
Answer #1

(a)

Let F be the full-time employee level, Hj, Lj, and Wj be the number of temporary workers hired, laid off, and present in the week-j for j=1,2,...,8. Assume W0=0.

Minimize Z = 8*610 F + 450 ∑j Wj

Subject to,

360 F + 270 Wj = Dj for j=1,2,...,8 Where Dj is number of orders in the j-th week

1.1 F + 2.7 Wj <= 200 for j=1,2,...,8

Wj = Wj-1 + Hj - Lj for j=1,2,...,8

(b)

Solution

Week Full-timers (F) Part-timers (Wj) Hire (Hj) Lay-off (Lj) Productions Orders Rejections Wj-1 +
Hj - Lj
0 0
1 54 1 1 0 19500 19500 61 1
2 54 6 6 0 21000 21000 76 6
3 54 23 17 0 25600 25600 122 23
4 54 29 6 0 27200 27200 138 29
5 54 52 23 0 33400 33400 200 52
6 54 39 0 13 29800 29800 164 39
7 54 29 0 10 27000 27000 136 29
8 54 43 15 0 31000 31000 176 43
Totals 429 223
Cost/ unit 610 450 Total cost
Cost 261568 100220 361788